# Performance of multiple filter-cavity schemes for frequency-dependent squeezing in gravitational-wave detectors

### v1

@@ V Conclusions and future steps
Frequency dependent squeezing has been proved an effective techniques to reduce quantum noise in gravitational-wave detectors. In this work we have extended the previous studies about FDS requiring complex cavity systems to achieve the full rotation of the squeezing ellipse, and applied the results to ET-LF. We have seen that, while the constraints on the CFC parameters are stringent, they are readily achievable and imperfections can be partially compensated. Importantly, the CFC configuration behaves equally or better in terms of robustness with respect to squeezing degradation sources compared to the 2FC one. In particular, we have provided new theoretical insights in understanding cascaded mode mismatch on the degradation of squeezed light. For future work, special attention should nevertheless be given to the realization of the middle mirror’s parameters, and controllability issues of the CFC scheme should be addressed for this precise configuration, expanding for example on [40]. We also note that our insights on quantum degradation in multiple cavity schemes could be extended to hybrid EPR+1FC [22] or quantum teleportation schemes [20], where the interferometer itself acts as an effective single or multi-cavity.
Armed with the detailed model of quantum noise, we have studied the tuned ET-LF + 1FC configuration – similar to the layout of LIGO during the O4 run – and compared it to the final detuned ET-LF + FDS design. This comparison is relevant because the footprint of both layouts are similar, a crucial point for an underground interferometer. The relatively modest redshift gain of the detuned configuration compared to the tuned case incites us to consider a 2-step implementation for ET: first attain the tuned configuration by removing the middle mirror of CFC, then, add the latter and detune the interferometer to achieve the designed ET-LF sensitivity. Such an approach, should its controllability be demonstrated, may simplify the overall optical scheme, and mitigate the effects of some squeezing degradation sources.
Acknowledgements
We thank Michel Lequime, Myriam Zerrad, Anne Daumas, Michael Hartmann, Giancarlo Cella, Jérôme Degallaix, and Matthew Evans for the fruitful discussions and comments. This work has been supported by ANR-23-CE31-0004, by the Paris Region and by the EU Horizon 2020 Research and Innovation Program under the Marie Sklodowska-Curie Grant Agreement No. 101003460 ”PROBES”.
@@ A.3 Quantum noise in passive systems
We assume that the input SDM is a squeezed rotated state:
$\bm{\bar{S}}^{\text{in}}=\bm{R}(\phi)\bm{S}(2r)\bm{R}(\phi)^{T}$ (32)
with $\bm{S}(r)=\text{diag}(e^{r},e^{-r})$ the squeezing matrix. Noting that $\bm{R}(\theta_{D})$ commutes with $\bm{\sigma}$ and that the latter matrix is hermitian, we find that
$\bm{\bar{S}}^{\text{out}}=\bm{R}(\theta_{D}+\phi)\left[(C\bm{1}-D\bm{\sigma})(\bm{S}(2r)-\bm{1})(C\bm{1}-D\bm{\sigma})+\bm{1}\right]\bm{R}(\theta_{D}+\phi)^{T}$ (33)
@@
where the homodyne readout is $u(\theta)=\begin{bmatrix}\cos(\theta)\\ \sin(\theta)\end{bmatrix}$, and $\Lambda_{\pm}^{\mathbb{R}}\coloneqq\eta\left[(1-\Xi)e^{\pm 2r}+\Xi e^{\mp 2r}\right]+(1-\eta)$ are the eigenvalues of $\text{Re}(\bm{\bar{S}^{\prime}})$. This can be seen by noting that it is a diagonal matrix per Eq. 35.
@@ A.4 Phase noise
Phase noise on a parameter $X_{k}$ on which the spectral density depends is accounted by effectively by averaging its expression in the interval $[X_{k}-\delta X_{k},X_{k}+\delta X_{k}]$:
$\begin{split}\bar{S}^{\text{avg}}=&\bar{S}(X_{1},\dots,X_{n})\\ &+\sum_{k}\left(\frac{\bar{S}(X_{k}+\delta X_{k})+\bar{S}(X_{k}-\delta X_{k})}{2}-\bar{S}(X_{k})\right)\\ \end{split}$ (41)
@@ Appendix B Proof of the scaling laws for bandwidth, detuning and loss
In this appendix we prove the equations Eqs. 1, 2, 4 and 3 as well as the observation of Section III.1, that the responses of 2FC and CFC are similar when their RTL $\Lambda$ and $\Lambda^{\prime}$ are equal. Our proof generalizes that of [22] so this latter observation.
@@ B.3 Coupled Filter Cavity with symmetric loss
We now expand the expressions of $\phi_{i}=-2(\Omega+\Delta\omega_{i})\tau$ for $i\in\{1,2,a,c\}$, while $\tau=L/c$ where $L$ is the length the cavity, assumed to be identical for all four cavities. We then identify both sides of Eqs. 50 and 51 by powers of $\Omega$ and assuming $T_{a}\ll T_{c}$:
@@
where the bandwidths $\gamma_{i}=\frac{cT_{i}}{4L}$ correspond to the ones in the lossless case. Thus, the scaling laws Eqs. 1, 2, 4 and 3 are obtained by solving the previous equations assuming $\Lambda^{\prime}=0$ and $\Lambda=0$. Then, we obtain in the lossy case that $\Lambda^{\prime}=\Lambda$ by looking at Eqs. 54 and 55, the left sides being zero. This concludes the proof of Eqs. 1, 2, 4 and 3 and formally justifies the observation of Section III.1.
@@ Appendix C Substrate loss of the middle mirror in CFC
In this appendix we justify the statement of Section III.1 that substrate loss of the middle mirror in CFC has a negligible impact on the degradation of squeezing. We consider a middle mirror made of fused silica and 5 cm thick. Substrate loss in Fused Silica core fibers are of the order of 1 ppm / cm [42, 43]. We take a conservative 2 ppm/cm figure for bulk fused silica, which amounts to an added round trip loss of 20 ppm. Using this value, and adding the 30 ppm of loss per subcavity (see Table 1), we compare the squeezing performances when the substrate is facing towards the c-cavity (with High Reflectivity coating faced towards the a-cavity) to when it is facing the a-cavity. From the results in Fig. 10, we see that by placing the substrate in the c-cavity, squeezing is never degraded by more than 0.5 dB compared to the 2FC scheme.
@@
Appendix D Mode mismatch in a coupled filter cavity
In this appendix we compute the mode mismatch between the two sub-cavities of a CFC setup. We consider the configuration relevant to this paper, that is, in the perfect case, both input and end mirrors (indices $c$ and $e$) are convex with the same radius of curvature $R_{a}=R_{c}=R$, are separated by a length $L$ such that $L<2R$ (for stability reasons) and the middle mirror (index $a$)is flat at the position $L/2$. This forms two plano-concave cavities, with identical waist position (at the position of the middle mirror) and waist size (since the input and end mirrors have the same RoC).
Now, for a realistic cavity, assume that the radii of curvatures are $R_{c}=R(1+\epsilon_{c})$, $\infty>R_{a}\gg R$ and $R_{e}=R(1+\epsilon_{e})$, with $\epsilon_{a,e}\ll 1$. The middle mirror is still placed at $L/2$.
Since we are interested in the relative mode mismatch between the two subcavities, we can assume without loss of generality that $\epsilon_{c}=0$ and concentrate all the defects on the parameter $\epsilon_{e}$ of the end mirror. This is represented in Fig. 11.
Figure 11: Realistic CFC layout with mode mismatch. The upper figure shows the CFC with perfect mirrors: the waist of the two subcavities is at the position of the middle mirror., and both curved mirrors have the same radius of curvature. The bottom figure shows imperfections in the a-cavity.
Under this assumption, the fundamental eigenmode of the first sub-cavity is the same as in the perfect case, while the second cavity mode is changed. To determine the mode mismatch (i.e. the mode overlap between the two fundamental eigenmodes), we first compute the position and the size of the waist of the second sub-cavity.
First, let $z=0$ be the middle of the cavity, at the position of the middle mirror. The waist position of the second cavity, $z^{\prime}_{0}$, is given by
$z^{\prime}_{0}=\frac{d(R(1+\epsilon_{e})-d)}{R_{a}+R(1+\epsilon_{e})-2d}$ (56)
where a positive RoC value corresponds to a concave mirror and $d=L/2$. The waist squared is
$w_{0}^{\prime 2}=\frac{\lambda}{\pi}\sqrt{z_{0}^{\prime}(R_{a}-z_{0}^{\prime})}$ (57)
In the perfect case, the waist position and size are
$z_{0}=0,\quad w_{0}^{2}=\frac{\lambda}{\pi}\sqrt{d(R-d)}$ (58)
In the imperfect case, at lowest order in $\epsilon$ and $\frac{1}{R_{a}}$, the waist position is
$z^{\prime}_{0}=\frac{d(R-d)}{R_{a}}$ (59)
and the waist size is
$w^{\prime}_{0}=w_{0}\left(1+\epsilon\frac{R}{4(R-d)}-\frac{R-2d}{4R_{a}}\right)$ (60)
The complex mismatch $m=\sqrt{\Upsilon_{c}}\,e^{\mathrm{i}\psi_{mmc}}$ of the cavity is [35]:
$m$ $=\frac{w^{\prime}_{0}-w_{0}}{w_{0}}+\mathrm{i}\frac{z_{0}^{\prime}}{2\sqrt{d(R-d)}}$ (61)
$=\epsilon\frac{R}{4(R-d)}-\frac{R-2d}{4R_{a}}+\mathrm{i}\frac{\sqrt{d(R-d)}}{2R_{a}}$ (62)
from which we identify $\Upsilon_{c}=|m|^{2}$ and $\psi_{mmc}=\arg(m)$.
Numerically, state of the art mirrors fabrication allows to obtain for a flat mirror $R_{a}>10^{9}$ m. This lower bound is estimated by making the assumption that the RMS flatness $\sigma\sim 10^{-10}$ m of a state-of-the art mirror is due to its sphericity. Then the radius of curvature is of the order of $D^{2}/\sigma$ where $D\sim 0.3$ m is the diameter of the mirror. In practice, since mirror maps show no quadratic contribution to $\sigma$ (see [24]), we have $R_{a}>D^{2}/\sigma$. We further have $R\epsilon\sim 1$ m, for typical $R=15$ km. We thus expect $\Upsilon_{c}\sim 10^{-9}$ for $L=1-10$ km. This shows that, provided state-of-the art mirror quality, the internal mode mismatch of a coupled cavity is negligible.
Likewise, if the length of the second cavity changes from $d$ to $d+\delta$ with $\delta\ll d$, the effect at lowest order is to change the waist size to be
$w_{0}^{\prime}=w_{0}+\left(1+\frac{\delta R}{4d(R-d)}\right)$ (63)
which means that the mode mismatch gets the contribution
$m=\frac{w^{\prime}_{0}-w_{0}}{w_{0}}=\frac{\delta R}{4d(R-d)}$ (64)
For $\delta\sim 1$ m and the same values for $R$ and $d$ as previously, we get $m\sim 10^{-4}$ so $\Upsilon_{a}\sim 10^{-8}$. Again, this shows that the mismatch due to cavity length defects is negligible.

### v2

@@ V Conclusions and future steps
Frequency dependent squeezing has been proved an effective techniques to reduce quantum noise in gravitational-wave detectors. In this work we have extended the previous studies about FDS requiring complex cavity systems to achieve the full rotation of the squeezing ellipse, and applied the results to ET-LF. We have seen that a 10-km CFC layout is achievable with standard mirror technologies, and deviations from nominal mirror parameters can be compensated. Importantly, the CFC configuration behaves equally or slightly better in terms of robustness with respect to squeezing degradation sources compared to the 2FC one. In particular, we have provided new theoretical insights in understanding the equivalency of CFC and 2FC layouts for round trip losses, as well as the propagation of cascaded mode mismatch on the degradation of squeezed light. For future work, special attention should nevertheless be given to controllability issues of the CFC scheme, expanding for example on [45]. We also note that our insights on quantum degradation in multiple cavity schemes could be extended to hybrid EPR+1FC [24] or quantum teleportation schemes [22], where the interferometer itself acts as an effective single or multi-cavity.
Armed with the detailed model of quantum noise, we have studied the tuned ET-LF + 1FC configuration – similar to the layout of LIGO during the O4 run – and compared it to the final detuned ET-LF + 2FC/CFC designs. This comparison is relevant because the footprint of both layouts are similar, a crucial point for an underground interferometer. The relative redshift gain of the detuned configuration compared to the tuned case incites us to consider a 2-step implementation for ET: first attain the tuned configuration by removing the middle mirror of CFC, then, add the latter and detune the interferometer to achieve the designed ET-LF sensitivity. Such an approach, should its controllability be demonstrated, may simplify the overall optical scheme, and mitigate the effects of some squeezing degradation sources. Finally, we note that, up to some overall geometrical factor, the results of this paper are independent of the final geometry of the detector, thus remain relevant should either a triangular or a L-shape configuration be ultimately adopted.
Acknowledgments
We thank Michel Lequime, Myriam Zerrad, Anne Daumas, Michael Hartmann, Giancarlo Cella, Jérôme Degallaix, Francesco Di Renzo, Mikhail Korobko, Léon Vidal and Matthew Evans for the fruitful discussions and comments. This work has been supported by Quantum FRESCO (ANR-23-CE31-0004), by the Paris Region and by the EU Horizon 2020 Research and Innovation Program under the Marie Sklodowska-Curie Grant Agreement No. 101003460 ”PROBES”.
@@ A.3 Quantum noise in passive systems
We assume that the input SDM is a squeezed rotated state $\bm{\bar{S}}^{\text{in}}=\bm{R}(\phi)\bm{S}(2r)\bm{R}(\phi)^{T}$, with $\bm{S}(r)=\text{diag}(e^{r},e^{-r})$ the squeezing matrix. Noting that $\bm{R}(\theta_{D})$ commutes with $\bm{\sigma}$ and that the latter matrix is hermitian, we find that
$\begin{split}\bm{\bar{S}}^{\text{out}}=\bm{R}(\theta_{D}+\phi)&\left[(C\bm{1}-D\bm{\sigma})(\bm{S}(2r)-\bm{1})(C\bm{1}-D\bm{\sigma})\right.\\ &\quad\left.+\bm{1}\right]\bm{R}(\theta_{D}+\phi)^{T}\end{split}$ (32)
@@
where the homodyne readout is $\bm{u}(\theta)=\begin{bmatrix}\cos(\theta)\\ \sin(\theta)\end{bmatrix}$, and $\Lambda_{\pm}^{\mathbb{R}}\coloneqq\eta\left[(1-\Xi)e^{\pm 2r}+\Xi e^{\mp 2r}\right]+(1-\eta)$ are the eigenvalues of $\text{Re}(\bm{\bar{S}^{\prime}})$. This can be seen by noting that it is a diagonal matrix per Eq. 34.
@@ A.4 Phase noise
Phase noise on a parameter $X_{k}$ on which the spectral density depends is accounted by effectively by averaging its expression in the interval $X_{k}-\delta X_{k},X_{k}+\delta X_{k}]$:
[…]
@@ Appendix B Proof of the scaling laws for bandwidth, detuning and loss
In this appendix we prove the equations Eqs. 1, 2, 4 and 3 as well as the observation of Section III.1, that the responses of 2FC and CFC are similar when their RTL $\Lambda$ and $\Lambda^{\prime}$ are equal. Our proof generalizes that of [24] to this latter observation.
@@ B.3 Coupled Filter Cavity with symmetric loss
We now expand the expressions of $\phi_{i}=-2(\Omega+\Delta\omega_{i})\tau$ for $i\in\{1,2,a,c\}$, while $\tau=L/c$ where $L$ is the length the cavity, assumed to be identical for all four cavities. We then identify both sides of Eqs. 49 and 50 by powers of $\Omega$ and assuming $T_{a}\ll T_{c}$:
@@
where the bandwidths $\gamma_{i}=\frac{cT_{i}}{4L}$ correspond to the ones in the lossless case. Thus, the scaling laws Eqs. 1, 2, 4 and 3 are obtained by solving the previous equations assuming $\Lambda^{\prime}=0$ and $\Lambda=0$. Then, we obtain in the lossy case that $\Lambda^{\prime}=\Lambda$ by looking at Eqs. 53 and 54, the left sides being zero. This concludes the proof of Eqs. 1, 2, 4 and 3 and formally justifies the observation of Section III.1.
@@ Appendix C Substrate loss of the middle mirror in CFC
In this appendix we justify the statement of Section III.1 that substrate loss of the middle mirror in CFC has a negligible impact on the degradation of squeezing. We consider a middle mirror made of fused silica and 5 cm thick. Substrate loss in bulk fused silica depend on the dryness of the substrate and can be less than 1 ppm / cm [47, 48]. We retain a conservative 2 ppm/cm figure for bulk fused silica, which amounts to an added round trip loss of 20 ppm. Using this value, and adding the 30 ppm of loss per subcavity (see Table 1), we compare the squeezing performances when the substrate is facing towards the c-cavity (with High Reflectivity coating faced towards the a-cavity) to when it is facing the a-cavity. From the results in Fig. 11, we see that by placing the substrate in the c-cavity (green curve), squeezing is never degraded by more than 0.5 dB compared to the 2FC scheme.
@@
Appendix D Geometrical defects in a coupled cavity
In this section, we derive the mode mismatch amplitude for a coupled filter cavity with geometric imperfections, to justify that with current state-of-the-art optics, the intra-cavity mode mismatch $\Upsilon_{a}$ of Table 1 can reasonably be neglected compared to the other mismatch contributions of this paper; in other words we show that $\Upsilon_{a}\ll 1\%$.
Figure 12: (a): Ideal coupled filter cavity: both input and end mirrors are concave with radius of curvature $R_{0}$; the middle mirror is plane and placed in the middle of the cavity, thus delimiting two sub-cavities of equal length. (b): Realistic coupled filter cavity, including radius error $\varepsilon_{e}$, middle mirror placement error $\Delta$, residual spheric interfaces $R_{m1}$, $R_{m2}$, and finite substrate thickness $t$ and index $n$ (the high-reflectivity surface is oriented towards the a-cavity).
D.1 Review of Gaussian optics
D.1.1 Gaussian beams
The $q$ parameter is defined for a Gaussian beam propagating along $z$ as $q(z)=(z-z_{0})+\mathrm{i}z_{R}$ where $z_{0}$ is the position of the waist and $z_{R}$ is the Rayleigh range. Equivalently, we write
$\frac{1}{q(z)}=\frac{1}{R(z)}-\frac{2\mathrm{i}}{kw(z)^{2}}$ (55)
where $R(z)=\frac{(z-z_{0})^{2}+z_{R}^{2}}{z-z_{0}}$ is the spherical phase curvature of the beam and $w(z)=\sqrt{\frac{2}{kz_{R}}}\sqrt{(z-z_{0})^{2}+z_{R}^{2}}$ is the waist of the beam at position $z$. We define a normalized gaussian mode $u(r,z)$ of the electric field as
$\begin{split}u(r,z)&=\sqrt{\frac{2}{\pi w(z)^{2}}}\exp\left[-\frac{\mathrm{i}kr^{2}}{2q(z)}\right]\exp\left[\mathrm{i}\varphi(z)\right]\\ &=\sqrt{-\frac{k}{\pi}\operatorname{Im}\left(\frac{1}{q(z)}\right)}\exp\left[-\frac{\mathrm{i}kr^{2}}{2q(z)}\right]\exp\left[\mathrm{i}\varphi(z)\right]\end{split}$ (56)
[…]
D.1.2 Overlap integral and mode mismatch amplitude
We are interested in the overlap integral between two modes $u_{1}$ and $u_{2}$ of parameters $q_{1}(z),q_{2}(z)$:
[…]
$\mathcal{O}(z)=\frac{2\mathrm{i}}{\frac{1}{q_{1}^{*}}-\frac{1}{q_{2}}}\sqrt{\operatorname{Im}\left(\frac{1}{q_{1}}\right)\operatorname{Im}\left(\frac{1}{q_{2}}\right)}\ e^{\mathrm{i}(\varphi_{2}-\varphi_{1})}$ (58)
where we omitted the dependency in $z$ for clarity. Using the relation $\operatorname{Im}\left(\frac{1}{q}\right)=\frac{-z_{R}}{|q|^{2}}$, this simplifies further to
$\mathcal{O}(z)=2\frac{\sqrt{z_{R1}z_{R2}}}{q_{2}-q_{1}^{*}}\frac{\mathrm{i}q_{1}^{*}q_{2}}{|q_{1}q_{2}|}e^{\mathrm{i}(\varphi_{2}-\varphi_{1})}$ (59)
Note that the second fraction has modulus 1, so since we are only interested in the amplitude of this overlap integral, we get
$|\mathcal{O}(z)|=2\frac{\sqrt{z_{R1}z_{R2}}}{|q_{2}-q_{1}^{*}|}$ (60)
which is indeed independent of $z$ since $q_{2}-q_{1}^{*}=z_{01}-z_{02}+\mathrm{i}(z_{R2}+z_{R1})$. Finally, the modulus square is related to the mode matching as
$1-\Upsilon\coloneqq|\mathcal{O}|^{2}=\frac{4z_{R1}z_{R2}}{|q_{2}-q_{1}^{*}|^{2}}$ (61)
It is possible to obtain a compact expression of the mode mismatch amplitude $\Upsilon$ using the above expression at lowest order in $\Delta z_{0}\coloneqq z_{01}-z_{02}$ and $\Delta z_{R}\coloneqq z_{R1}-z_{R2}$. Since $|q_{2}-q_{1}^{*}|^{2}\simeq\Delta z_{0}^{2}+4z_{R}^{2}\simeq 4z_{R}^{2}$, we have
$\Upsilon\simeq\frac{\Delta z_{0}^{2}}{4z_{R}^{2}}+\frac{\Delta z_{R}^{2}}{4z_{R}^{2}}$ (62)
Equivalently, in terms of waists, we use $z_{Ri}=kw_{0i}^{2}/2$ to obtain $\Upsilon={\frac{\Delta z_{0}^{2}}{4z_{R}^{2}}+\frac{\Delta w_{0}^{2}}{w_{0}^{2}}}$, where $\Delta w_{0}=w_{01}-w_{02}$.
D.1.3 Fundamental gaussian mode in a linear cavity
In the rest of this appendix, the longitudinal coordinate grows left to right. For a two-mirror cavity of length $L$ with input/end mirror radii of curvature $R_{i},R_{e}$ (using the convention that, seen from inside the cavity, a mirror has positive radius of curvature if it is concave), the waist position $z_{0}$ measured from the first mirror ($z=0$ at the mirror $R_{i}$), and its Rayleigh length squared $z_{R}^{2}$ are
$z_{0}$ $=\frac{L^{2}-R_{e}L}{\,2L-R_{i}-R_{e}\,}$, (63)
$z_{R}^{2}$ $=\frac{L(R_{i}-L)(R_{e}-L)(R_{i}+R_{e}-L)}{(R_{e}+R_{i}-2L)^{2}}.$ (64)
Note that for a perfect, flat–curved half-cavity (first mirror plane, second mirror $R_{0}$, length $L_{0}$), Eqs. 63 and 64 give $z_{0}^{(0)}=0$ and $z_{R}^{2}=L_{0}(R_{0}-L_{0})$.
D.2 Mode mismatch due to geometrical imperfections
Consider first the ideal situation shown in Fig. 12 (a), of a symmetric Fabry–Perot cavity of length $2L_{0}$ with end mirrors of the same radius $R_{0}$ and a plane mirror in the middle. Each sub-cavity (“c-cavity” and “a-cavity”) is stable with perfect fundamental Gaussian having its waist at the plane mirror and
$z_{R}=\sqrt{L_{0}\,(R_{0}-L_{0})}.$ (65)
We now add realistic imperfections, as shown in Fig. 12 (b):
  • (i)
    Right end-mirror curvature deviation: $R_{e}=R_{0}(1+\varepsilon_{e})$, with $|\varepsilon_{e}|\ll 1$;
  • (ii)
    Middle plane displaced: left gap $L_{1}=L_{0}+\Delta$, right gap $L_{2}=L_{0}-\Delta$, with $|\Delta|\ll L_{0},R_{0}$;
  • (iii)
    Middle optic: entrance surface curvature $R_{m1}$, back (HR-coated) surface curvature $R_{m2}$, both large in magnitude compared to all other radii of curvature;
  • (iv)
    Middle substrate: thickness $t\ll\Delta,L_{0},R_{0}$, refractive index $n>1$; the HR coating is on the back surface, so the left half-beam traverses the glass to reflect.
Note that since we are only interested in the relative mode mismatch between the two eigenmodes of the two sub-cavities, we have assumed without loss of generality that the input mirror is undeformed. Additionally, the impact of coating non-uniformity can be neglected compared to these four imperfections, as the coatings are much thinner than any other length scale above.
Label by $c$ (left) and $a$ (right) the fundamental Gaussian eigenmodes of the two imperfect half-cavities. We compute, to the lowest non-trivial order in the small quantities $\varepsilon_{e},\Delta,\,1/R_{m1},\,1/R_{m2},\,t$, the mode mismatch between the two eigenmodes. We will repeatedly use the geometry factors
$\begin{split}c_{\Delta}&:=\frac{L_{0}-\tfrac{R_{0}}{2}}{z_{R}},\\ c_{k}&:=\frac{z_{R}}{2}\,(2L_{0}-R_{0}),\\ c_{\varepsilon}&:=\frac{R_{0}\,z_{R}}{2\,(R_{0}-L_{0})}.\end{split}$ (66)
D.2.1 a-cavity
This cavity is bounded by the middle HR surface (first mirror) and the right end mirror, so $L=L_{0}-\Delta$, $R_{i}=R_{m2}$ (large) and $R_{e}=R_{0}(1+\varepsilon_{e})$.
Evaluating (63) with large $R_{i}$:
$z_{0a}=-\,\frac{L^{2}-R_{e}L}{R_{i}}=\frac{z_{R}^{2}}{R_{m2}}$ (67)
where we neglected terms of order $R_{i}^{-2}$. This is the position of the waist relative to the position of the HR surface of the middle mirror, itself displaced by $\Delta$ from the midpoint of the CFC. Thus the waist position relative to the midpoint of the CFC is
$\tilde{z}_{0a}=\,z_{0a}+\Delta=\Delta+\frac{z_{R}^{2}}{R_{m2}}$ (68)
Similarly, (64) reduces to lowest order in $R_{i}^{-1}$ to
$z_{Ra}^{2}$ $=L(R_{e}-L)\;-\;\frac{L(R_{e}-L)(R_{e}-2L)}{R_{i}}$ (69)
$=z_{R}^{2}-(R_{0}-2L_{0})\Delta+L_{0}R_{0}\varepsilon_{e}+z_{R}^{2}\frac{2L_{0}-R_{0}}{R_{m2}}$ (70)
This means that the Rayleigh length is
$z_{Ra}=z_{R}+c_{\Delta}\,\Delta+c_{\varepsilon}\,\varepsilon_{e}+\frac{c_{k}}{R_{m2}}.$ (71)
D.2.2 c-cavity
The left cavity is bounded by the input mirror (first mirror $R_{0}$) and the effective reflector seen in vacuum when the beam enters the substrate, reflects from the HR on the back surface, and exits. The effect of the substrate is to add optical path and change the radius of curvature of the high-reflective surface, such that, to first order, $L=L_{0}+\Delta+\frac{t}{n}=L_{0}+\tilde{\Delta}$, $R_{i}=R_{0}$ and $R_{e}=\frac{-R_{m2}}{n}$ (the minus sign is because $R_{m2}$ corresponds to the radius of curvature as seen from the a-cavity, outside the c-cavity). Note that $R_{m1}$ has no contribution at lowest order because the beam crosses from vacuum to substrate then back.
Expanding (63) for large $|R_{e}|$:
$z_{0c}=L+\frac{L(L-R_{i})}{R_{e}}=L_{0}+\tilde{\Delta}+\frac{nz_{R}^{2}}{R_{m2}}.$ (72)
This is the position of the waist relative to the input mirror, which is $L_{0}$ to the left of the midpoint of the CFC. This means that, relative to the midpoint, the waist is at position
$\tilde{z}_{0c}=z_{0c}-L_{0}=\Delta+\frac{t}{n}+\frac{nz_{R}^{2}}{R_{m2}}$ (73)
For the Rayleigh range, we use (64) with large $|R_{e}|$ to get:
$z_{Rc}^{2}=L(R_{i}-L)\;-\;\frac{L(R_{i}-L)(R_{i}-2L)}{R_{e}}.$ (74)
With $L=L_{0}+\tilde{\Delta}$ and $R_{i}=R_{0}$, we obtain
$z_{Rc}=z_{R}-c_{\Delta}\left(\Delta+\frac{t}{n}\right)-\frac{nc_{k}}{R_{m2}}$ (75)
D.2.3 Application
[…]